OpenStudy (anonymous):

a client comes to you for an investment advice on his $500,000 winnings from the lottery. which of these options would be the best for his investment 6% compounded interest quarterly for 5 years or 8% compounded interest annually for 5 years or 14.5% simple interest for 10 years. Please explain how you coome up with your answer

OpenStudy (anonymous):

Call your options a,b, and c as in order. a) 500000(1.015)^20 = 673427.50. b) 500000(1.08)^5 = 734664.04. c) 500000 + (10)(500000)(0.145) = 1225000. Simple interest gives the most in this example.